OpenStudy (anonymous):

Expand the logarithmic expression logb square root of 13/74 1. 1/2 logb 13 - 1/2logb 73 2. 1/2 logb 13 + 1/2logb 73 3. logb (13-73) 4. sqrt logb (13-73)

OpenStudy (anonymous):

\[\log_b(\sqrt{\frac{13}{17}})\]?

OpenStudy (anonymous):

no 13/74

OpenStudy (anonymous):

\[\log_b\left(\sqrt{\frac{13}{73}}\right)\] like that?

OpenStudy (anonymous):

haha yes that's it

OpenStudy (anonymous):

first rewrite the radical in exponential notation as \[\log_b\left({\frac{13}{73}}\right)^{\frac{1}{2}}\] then the \(\frac{1}{2}\) comes out front as \[\frac{1}{2}\log_b(\frac{13}{73})\]

OpenStudy (anonymous):

after that use \(\log(\frac{A}{B})=\log(A)-\log(B)\) to give you \[\frac{1}{2}\left(\log_b(13)-\log_b(73)\right)\]

OpenStudy (anonymous):

and finally distribute the \(\frac{1}{2}\)

OpenStudy (anonymous):

so, 1/2 logb 13- 1/2logb 73?